# Coarse-to-fine Q-Network with Action Sequence for Data-Efficient Robot Learning

### v5

@@ 4 Experiment
We study CQN-AS on 25 humanoid control tasks from BiGym (Chernyadev et al., 2024) and 20 tabletop manipulation tasks from RLBench (James et al., 2020) (see Figure 4 for examples of robotic tasks). To focus on challenging robotic tasks that aim to induce policies generating realistic behaviors, we consider a practical setup of demo-driven RL where we initialize training with a modest amount of expert demonstrations and then train with online data. In particular, our experiments are designed to investigate the following questions:
@@
    Can CQN-AS quickly match the performance of a recent BC algorithm (Zhao et al., 2023) and surpass it through online learning? How does CQN-AS compare to previous model-free RL algorithms (Yarats et al., 2022; Seo et al., 2024)?
@@
We consider model-free RL baselines that learn deterministic policies, as we find that stochastic policies struggle to solve fine-grained manipulation tasks. Specifically, we consider (i) Coarse-to-fine Q-Network (CQN; Seo et al. 2024), our backbone algorithm and (ii) DrQ-v2+, an optimized demo-driven variant of an actor-critic algorithm DrQ-v2 (Yarats et al., 2022) that uses a deterministic policy algorithm and data augmentation. We further consider (iii) Action Chunking Transformer (ACT; Zhao et al. 2023) that trains a transformer (Vaswani et al., 2017) policy to predict action sequences from expert demonstrations and utilizes temporal ensemble, as our BC baseline.
@@
For training with expert demonstrations, we follow the setup of Seo et al. (2024). We keep a separate replay buffer that stores demonstrations and sample half of training data from demonstrations. We also relabel successful online episodes as demonstrations and store them in the demonstration replay buffer. For CQN-AS, we use an auxiliary BC loss from Seo et al. (2024) based on large margin loss (Hester et al., 2018). For actor-critic baselines, we use an auxiliary BC loss that minimizes L2 loss between the policy outputs and expert actions.
Figure 5: BiGym results on 25 sparsely-rewarded mobile bi-manual manipulation tasks. All experiments are initialized with 17 to 60 human-collected demonstrations, and RL methods are trained with an auxiliary BC objective. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 8 runs.
@@ 4.1 BiGym Experiments
We study CQN-AS on mobile bi-manual manipulation tasks from BiGym (Chernyadev et al., 2024). BiGym’s human-collected demonstrations are often noisy and multi-modal, posing challenges for RL algorithms. These algorithms must effectively leverage the information within demonstrations to learn strong initial behaviors, thereby mitigating exploration difficulties in sparsely rewarded tasks.
@@
We consider 25 BiGym tasks with 17 to 60 demonstrations. We use RGB observations with 84$\times$84 resolution from head, left_wrist, and right_wrist cameras. We also use low-dimensional proprioceptive states. We use (i) absolute joint position control action mode and (ii) floating base that replaces locomotion with classic controllers. We use the same set of hyperparameters for all the tasks. Details on BiGym experiments are available in Appendix B.
@@
Figure 5 shows that CQN-AS quickly matches the performance of ACT and outperforms it through online learning on most tasks, while other RL algorithms fail to do so especially on challenging long-horizon tasks such as Move Plate and Saucepan To Hob. A notable result here is that CQN-AS enables solving challenging BiGym tasks while other RL baselines completely fail as they achieve a 0% success rate on many tasks.
@@
However, CQN-AS struggles to achieve meaningful success rate on some of the long-horizon tasks that require interaction with delicate objects such as cups or cutlery. This leaves room for future work to incorporate advanced vision encoders (He et al., 2016) or critic architectures (Chebotar et al., 2023; Springenberg et al., 2024).
Figure 6: RLBench results on 20 sparsely-rewarded tabletop manipulation tasks from RLBench (James et al., 2020). All experiments are initialized with 100 synthetic demonstrations generated via motion-planning and RL methods are trained with an auxiliary BC objective. As expected, with synthetic demonstrations, CQN-AS achieves similar performance to CQN on most tasks. However, CQN-AS often significantly outperforms baselines on several challenging, long-horizon tasks such as Open Oven. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
@@ 4.2 RLBench Experiments
We also study CQN-AS on manipulation tasks from RLBench (James et al., 2020). Unlike BiGym, RLBench provides synthetic demonstrations generated via motion planning, which are cleaner and more consistent. This allows us to examine whether CQN-AS is also effective in settings with clean, unambiguous demonstrations – where the effect of each single-step action is easier to interpret.
@@
We use the official CQN implementation for collecting demonstrations and reproducing the baseline results on the same set of tasks. We use RGB observations with 84$\times$84 resolution from front, wrist, left_shoulder, and right_shoulder cameras. We also use low-dimensional proprioceptive states consisting of 7-dimensional joint positions and a binary value for gripper open. We use 100 demonstrations and delta joint position control action mode. We use the same set of hyperparameters for all the tasks, in particular, we use action sequence of length 4. More details on RLBench experiments are available in Appendix B.
@@
Because RLBench provides synthetic clean demonstrations, as we expected, Figure 6 shows that CQN-AS achieves similar performance to CQN on most tasks, except 2/25 tasks where it hurts the performance. But we still find that CQN-AS achieves quite superior performance to CQN on some challenging long-horizon tasks such as Open Oven or Take Plate Off Colored Dish Rack. These results show that CQN-AS can be used in various benchmark with different characteristics.
@@
(d) Effect of temporal ensemble
@@
Figure 7: Ablation studies and analysis on the effect of (a) action sequence, (b) RL objective, (c) $N$-step return, and (d & e) temporal ensemble. (f) We also provide results on locomotion tasks from DeepMind Control Suite (Tassa et al., 2020), where CQN-AS fails to improve performance. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
@@ 4.3 HumanoidBench Experiments
To show that CQN-AS is generally applicable to tasks without demonstrations, we also study CQN-AS on densely-rewarded tasks from HumanoidBench (Sferrazza et al., 2024). Due to space constraints, we describe the setup and results in Appendix A. In summary, we find that CQN-AS significantly improves the performance of base underlying algorithm CQN, also outperforming a model-free RL baseline, SAC (Haarnoja et al., 2018). This shows the general applicability of our idea.
@@ 4.4 Ablation Studies, Analysis, Failure Cases
Figure 7a shows the performance of CQN-AS with different action sequence lengths. We find that training the critic network with longer action sequences tends to consistently improve performance.
@@
Figure 7b shows the performance of CQN-AS without RL objective that trains the model only with BC objective on successful demonstrations. We find this baseline significantly underperforms CQN-AS, which shows that RL objective enables the agent to learn from trial-and-error experiences.
@@
Figure 7c shows experimental results with varying $N$-step returns. We find that too high $N$-step return significantly degrades performance. We hypothesize this is because the variance from $N$-step return makes it difficult to learn useful value functions.
@@
Figure 7d shows that performance degrades without temporal ensemble on Saucepan To Hob as temporal ensemble induces a smooth motion and thus improves performance in fine-grained control tasks. But we also find that temporal ensemble is harmful on Reach Target Single. This is because temporal ensemble uses predictions from previous steps and thus makes it difficult to refine behaviors based on recent observations. Nonetheless, we use temporal ensemble for all the tasks as it helps on most tasks and we aim to use the same set of hyperparameters.
@@
We further provide results with different temporal ensemble magnitudes by adjusting a hyperparameter $m$ in Figure 7e. Here, higher $m$ puts higher weights on recent actions and thus very high $m$ corresponds to using only first action. Similarly to previous paragraph, we find that higher $m$ leads to better performance on Reach Target Single that needs fast reaction, but degrades performance on Saucepan To Hob.
@@
Figure 7f shows that CQN-AS underperforms CQN on locomotion tasks with torque control, which are drawn from the DeepMind Control Suite (Tassa et al., 2020). We hypothesize that this performance degradation arises because sequences of joint positions tend to have clearer semantic structure in joint space, making them easier to learn from compared to sequences of raw torques. Addressing this failure case represents an interesting direction for future work.
@@ 5 Related Work
Recent behavior cloning approaches have shown that predicting a sequence of actions enables the policy to imitate noisy expert trajectories and helps in dealing with idle actions from human pauses during data collection (Zhao et al., 2023; Chi et al., 2023). Notably, Zhao et al. (2023) train a transformer model (Vaswani et al., 2017) that predicts action sequence and Chi et al. (2023) train a denoising diffusion model (Ho et al., 2020) that approximates the action distributions. This idea has been extended to multi-task setup (Bharadhwaj et al., 2024), mobile manipulation (Fu et al., 2024b) and humanoid control (Fu et al., 2024a). Our work is inspired by this line of research and proposes to learn RL agents with action sequence.
@@
In the context of RL, Medini & Shrivastava (2019) propose to pre-compute frequent action sequences from expert demonstrations and augment the action space with these sequences. However, this idea introduces additional complexity and is not scalable to setups without demonstrations. One recent work relevant to ours is Saanum et al. (2024) that encourage a sequence of actions from RL agents to be predictable and smooth. Our work differs in that we directly incorporate action sequences into value learning. Recently, Ankile et al. (2024) point out that RL with action sequence is challenging and instead propose to use RL for learning a single-step policy that corrects action sequence predictions from BC. In contrast, we show that RL with action sequence is feasible and improves performance of RL algorithms.
@@
Recent large language models have incorporated a notably similar idea to predicting action sequences from demonstrations – predicting multiple future tokens at once, or multi-token prediction (Gloeckle et al., 2024; Liu et al., 2024). For instance, Gloeckle et al. (2024) show that predicting multiple $n$ future tokens in parallel with $n$ independent output heads improves the performance and can accelerate inference speed. DeepSeek-V3 (Liu et al., 2024) also make a similar observation but with a sequential multi-token prediction. It would be interesting to see whether our idea can be utilized for fine-tuning these models with multi-token prediction.
@@
Approaches that learn RL agents with temporally extended high-level actions, or options, have been well studied (Sutton, 1988). The key idea is to train high-level policies that output options by manually defining subgoals (Kulkarni et al., 2016; Dayan & Hinton, 1992) or learning options from data (Bacon et al., 2017; Vezhnevets et al., 2017; Nachum et al., 2018), and then train a low-level agent that learns to execute low-level actions conditioned on options. Our work is not directly comparable to these works as we do not abstract temporally extended actions but use raw action sequences for value learning.
@@ 6 Discussion
One limitation of our work is the lack of real-world robot evaluation, as RL with fragile humanoid hardware remains a significant challenge. We hope our work serves as an important stepping stone toward addressing this issue. As discussed in Section 4.1 and Section 4.4, solving tasks involving small objects or requiring fine torque control also remains a limitation of our approach. We are excited about future directions, including real-world RL with humanoid robots, incorporating advanced critic architectures (Kapturowski et al., 2023; Chebotar et al., 2023; Springenberg et al., 2024), bootstrapping RL agents from imitation learning (Hu et al., 2023; Xing et al., 2024) or offline RL (Nair et al., 2020; Lee et al., 2021), extending the idea to recent model-based RL approaches (Hafner et al., 2023; Hansen et al., 2024), fine-tuning vision-language-action models that use action sequence (Team et al., 2024; Doshi et al., 2024) or language models that use multi-token prediction (Gloeckle et al., 2024; Liu et al., 2024) with our algorithm, to name but a few.
@@ Acknowledgements
Appendix A HumanoidBench Experiments
Setup
We follow a standard setup that trains RL agents from scratch. We use low-dimensional states consisting of proprioception and privileged task information as inputs. For tasks, we simply select the first 8 locomotion tasks in the benchmark. For baselines, we consider CQN and Soft Actor-Critic (SAC) (Haarnoja et al., 2018). For SAC, we use the results available from HumanoidBench repository, which are evaluated on tasks with dexterous hands. For CQN-AS and CQN, we also evaluate them on tasks with hands. We use the same set of hyperparameters for all the tasks (see Appendix B).
Comparison to baselines
Figure 8 shows that, by learning the critic network with action sequence, CQN-AS outperforms other model-free RL baselines, i.e., CQN and SAC, on most tasks. In particular, the difference between CQN-AS and baselines becomes larger as the task gets more difficult, e.g., baselines fail to achieve high episode return on Walk and Run tasks but CQN-AS achieves strong performance. This result shows that our idea of using action sequence can be applicable to generic setup without demonstrations.
Figure 8: HumanoidBench results on eight densely-rewarded humanoid control tasks (Sferrazza et al., 2024). All the experiments start from scratch and all the methods do not have an auxiliary BC objective. CQN-AS significantly improves the performance of underlying RL algorithm CQN, while outperforming a model-free RL baseline, SAC. For CQN-AS and CQN, we report the results aggregated over 8 runs. For SAC, we report the results aggregated over 3 runs available from public website. The solid line and shaded regions represent the mean and confidence intervals.

### v6

@@ 4 Experiment
We study CQN-AS on 25 humanoid control tasks from BiGym (chernyadev2024bigym) and 20 tabletop manipulation tasks from RLBench (james2020rlbench) (see Figure˜4 for examples of robotic tasks). To focus on challenging robotic tasks that aim to induce policies generating realistic behaviors, we consider a practical setup of demo-driven RL where we initialize training with a modest amount of expert demonstrations and then train with online data. In particular, our experiments are designed to investigate the following questions:
@@
    Can CQN-AS quickly match the performance of a recent BC algorithm (zhao2023learning) and surpass it through online learning? How does CQN-AS compare to previous model-free RL algorithms (yarats2022mastering; seo2024continuous)?
@@
We consider model-free RL baselines that learn deterministic policies, as we find that stochastic policies struggle to solve fine-grained manipulation tasks. Specifically, we consider (i) Coarse-to-fine Q-Network (CQN; seo2024continuous), our backbone algorithm and (ii) DrQ-v2+, an optimized demo-driven variant of an actor-critic algorithm DrQ-v2 (yarats2022mastering) that uses a deterministic policy algorithm and data augmentation. We further consider (iii) Action Chunking Transformer (ACT; zhao2023learning) that trains a transformer (vaswani2017attention) policy to predict action sequences from expert demonstrations and utilizes temporal ensemble, as our BC baseline.
@@
For training with expert demonstrations, we follow the setup of seo2024continuous. We keep a separate replay buffer that stores demonstrations and sample half of training data from demonstrations. We also relabel successful online episodes as demonstrations and store them in the demonstration replay buffer. For CQN-AS, we use an auxiliary BC loss from seo2024continuous based on large margin loss (hester2018deep). For actor-critic baselines, we use an auxiliary BC loss that minimizes L2 loss between the policy outputs and expert actions.
Figure 5: BiGym results on 25 sparsely-rewarded mobile bi-manual manipulation tasks. All RL algorithms are trained from scratch, with a replay buffer initialized with 17 to 60 human-collected demonstrations, and with an auxiliary BC objective. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 8 runs.
@@ 4.1 BiGym Experiments
We study CQN-AS on mobile bi-manual manipulation tasks from BiGym (chernyadev2024bigym). BiGym’s human-collected demonstrations are often noisy and multi-modal, posing challenges for RL algorithms. These algorithms must effectively leverage the information within demonstrations to learn strong initial behaviors, thereby mitigating exploration difficulties in sparsely rewarded tasks.
@@
We consider 25 BiGym tasks with 17 to 60 demonstrations [^2] BiGym benchmark provides different number of successful demonstrations for each task. But we use the same number of demonstrations for all algorithms. See Appendix B for more details.. We use RGB observations with $\times84$ resolution from head, left_wrist, and right_wrist cameras. We also use low-dimensional proprioceptive states. We use (i) absolute joint position control action mode and (ii) floating base that replaces locomotion with classic controllers. We use the same set of hyperparameters for all the tasks. Details on BiGym experiments are available in Appendix˜B.
@@
Figure˜5 shows that CQN-AS quickly matches the performance of ACT and outperforms it through online learning on most tasks, while other RL algorithms fail to do so especially on challenging long-horizon tasks such as Move Plate and Saucepan To Hob. A notable result here is that CQN-AS enables solving challenging BiGym tasks while other RL baselines completely fail as they achieve a 0% success rate on many tasks.
@@
However, CQN-AS struggles to achieve meaningful success rate on some of the long-horizon tasks that require interaction with delicate objects such as cups or cutlery. This leaves room for future work to incorporate advanced vision encoders (he2016deep) or critic architectures (chebotar2023q; springenberg2024offline).
Figure 6: RLBench results on 20 sparsely-rewarded tabletop manipulation tasks from RLBench (james2020rlbench). All RL algorithms are trained from scratch, with a replay buffer initialized with 100 synthetic demonstrations generated via motion-planning, and with an auxiliary BC objective. As expected, with synthetic demonstrations, CQN-AS achieves similar performance to CQN on most tasks. However, CQN-AS often significantly outperforms baselines on several challenging, long-horizon tasks such as Open Oven. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
@@ 4.2 RLBench Experiments
We also study CQN-AS on manipulation tasks from RLBench (james2020rlbench). Unlike BiGym, RLBench provides synthetic demonstrations generated via motion planning, which are cleaner and more consistent. This allows us to examine whether CQN-AS is also effective in settings with clean, unambiguous demonstrations – where the effect of each single-step action is easier to interpret.
@@
We use the official CQN implementation for collecting demonstrations and reproducing the baseline results on the same set of tasks. We use RGB observations with $\times84$ resolution from front, wrist, left_shoulder, and right_shoulder cameras. We also use low-dimensional proprioceptive states consisting of 7-dimensional joint positions and a binary value for gripper open. We use 100 demonstrations and delta joint position control action mode. We use the same set of hyperparameters for all the tasks, in particular, we use action sequence of length 4. More details on RLBench experiments are available in Appendix˜B.
@@
Because RLBench provides synthetic clean demonstrations, as we expected, Figure˜6 shows that CQN-AS achieves similar performance to CQN on most tasks, except 2/25 tasks where it hurts the performance. But we still find that CQN-AS achieves quite superior performance to CQN on some challenging long-horizon tasks such as Open Oven or Take Plate Off Colored Dish Rack. These results show that CQN-AS can be used in various benchmark with different characteristics.
Figure 7: HumanoidBench results on eight densely-rewarded humanoid control tasks (sferrazza2024humanoidbench). All the experiments start from scratch and all the methods do not have an auxiliary BC objective. CQN-AS significantly improves the performance of underlying RL algorithm CQN, while outperforming a model-free RL baseline, SAC. For CQN-AS and CQN, we report the results aggregated over 8 runs. For SAC, we report the results aggregated over 3 runs available from public website. The solid line and shaded regions represent the mean and confidence intervals.
@@
(d) Effect of temporal ensemble ($m=0.01$)
@@
Figure 8: Ablation studies and analysis on the effect of (a) action sequence, (b) RL objective, (c) NN-step return, and (d & e) temporal ensemble. (f) We also provide results on locomotion tasks from DeepMind Control Suite (tassa2020dm_control), where CQN-AS fails to improve performance. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
@@ 4.3 HumanoidBench Experiments
To show that CQN-AS is generally applicable to tasks without demonstrations, we also study CQN-AS on densely-rewarded tasks from HumanoidBench (sferrazza2024humanoidbench).
Setup
We follow a standard setup that trains RL agents from scratch. We use low-dimensional states consisting of proprioception and privileged task information as inputs. For tasks, we simply select the first 8 locomotion tasks in the benchmark. For baselines, we consider CQN and Soft Actor-Critic (SAC) (haarnoja2018soft). For SAC, we use the results available from HumanoidBench repository, which are evaluated on tasks with dexterous hands. For CQN-AS and CQN, we also evaluate them on tasks with hands. We use the same set of hyperparameters for all the tasks (see Appendix˜B).
Comparison to baselines
Figure˜7 shows that, by learning the critic network with action sequence, CQN-AS outperforms other model-free RL baselines, i.e., CQN and SAC, on most tasks. In particular, the difference between CQN-AS and baselines becomes larger as the task gets more difficult, e.g., baselines fail to achieve high episode return on Walk and Run tasks but CQN-AS achieves strong performance. This result shows that our idea of using action sequence can be applicable to generic setup without demonstrations.
@@ 4.4 Ablation Studies, Analysis, Failure Cases
Figure˜8a shows the performance of CQN-AS with different action sequence lengths. We find that training the critic network with longer action sequences tends to consistently improve performance, plateaus or decreases performance if the sequences get too long.
@@
Figure˜8b shows the performance of CQN-AS without RL objective that trains the model only with BC objective on successful demonstrations. We find this baseline significantly underperforms CQN-AS, which shows that RL objective enables the agent to learn from trial-and-error experiences.
@@
Figure˜8c shows experimental results with varying $N$-step returns. We find that too high $N$-step return significantly degrades performance. We hypothesize this is because the variance from $N$-step return makes it difficult to learn useful value functions.
@@
Figure˜8d shows that performance degrades without temporal ensemble on Saucepan To Hob as temporal ensemble induces a smooth motion and thus improves performance in fine-grained control tasks. But we also find that temporal ensemble is harmful on Reach Target Single. This is because temporal ensemble uses predictions from previous steps and thus makes it difficult to refine behaviors based on recent observations. Nonetheless, we use temporal ensemble for all the tasks as it helps on most tasks and we aim to use the same set of hyperparameters.
@@
We further provide results with different temporal ensemble magnitudes by adjusting a hyperparameter $m$ in Figure˜8e. Here, higher $m$ puts higher weights on recent actions and thus very high $m$ corresponds to using only first action. Similarly to previous paragraph, we find that higher $m$ leads to better performance on Reach Target Single that needs fast reaction, but degrades performance on Saucepan To Hob.
@@
Figure˜8f shows that CQN-AS underperforms CQN on locomotion tasks with torque control, which are drawn from the DeepMind Control Suite (tassa2020dm_control). We hypothesize that this performance degradation arises because sequences of joint positions tend to have clearer semantic structure in joint space, making them easier to learn from compared to sequences of raw torques. Addressing this failure case represents an interesting direction for future work.
@@ 5 Related Work
Recent behavior cloning approaches have shown that predicting a sequence of actions enables the policy to imitate noisy expert trajectories and helps in dealing with idle actions from human pauses during data collection (zhao2023learning; chi2023diffusion). Notably, zhao2023learning train a transformer model (vaswani2017attention) that predicts action sequence and chi2023diffusion train a denoising diffusion model (ho2020denoising) that approximates the action distributions. This idea has been extended to multi-task setup (bharadhwaj2024roboagent), mobile manipulation (fu2024mobile) and humanoid control (fu2024humanplus). Our work is inspired by this line of research and proposes to learn RL agents with action sequence.
@@
In the context of RL, medini2019mimicking propose to pre-compute frequent action sequences from expert demonstrations and augment the action space with these sequences. However, this idea introduces additional complexity and is not scalable to setups without demonstrations. One recent work relevant to ours is saanum2024reinforcement that encourage a sequence of actions from RL agents to be predictable and smooth. Our work differs in that we directly incorporate action sequences into value learning. Recently, ankile2024imitation point out that RL with action sequence is challenging and instead propose to use RL for learning a single-step policy that corrects action sequence predictions from BC. In contrast, we show that RL with action sequence is feasible and improves performance of RL algorithms.
@@
Recent large language models have incorporated a notably similar idea to predicting action sequences from demonstrations – predicting multiple future tokens at once, or multi-token prediction (gloeckle2024better; liu2024deepseek). For instance, gloeckle2024better show that predicting multiple $n$ future tokens in parallel with $n$ independent output heads improves the performance and can accelerate inference speed. DeepSeek-V3 (liu2024deepseek) also make a similar observation but with a sequential multi-token prediction. It would be interesting to see whether our idea can be utilized for fine-tuning these models with multi-token prediction.
@@
Approaches that learn RL agents with temporally extended high-level actions, or options, have been well studied (sutton1988learning). The key idea is to train high-level policies that output options by manually defining subgoals (kulkarni2016hierarchical; dayan1992feudal) or learning options from data (bacon2017option; vezhnevets2017feudal; nachum2018data), and then train a low-level agent that learns to execute low-level actions conditioned on options. Our work is not directly comparable to these works as we do not abstract temporally extended actions but use raw action sequences for value learning.
@@ 6 Discussion
One limitation of our work is the lack of real-world robot evaluation. Moreover, as discussed in Section˜4.1 and Section˜4.4, solving tasks involving small objects remains a limitation of our approach. One potential approach would be using strong pre-trained vision encoders, but we find that computational cost is often prohibitively large, which remains as an open problem. We are excited about future directions, including real-world RL with humanoid robots, incorporating advanced critic architectures (kapturowski2022human; chebotar2023q; springenberg2024offline), bootstrapping RL agents from imitation learning (hu2023imitation; xing2024bootstrapping) or offline RL (nair2020awac; lee2021offline), extending the idea to recent model-based RL approaches (hafner2023mastering; hansen2023td), extend parallel value learning scheme to autoregressive, multi-step Q-learning scheme (kahn2018self), fine-tuning vision-language-action models that use action sequence (team2024octo; doshi2024scaling) or language models that use multi-token prediction (gloeckle2024better; liu2024deepseek) with our algorithm, to name but a few.
@@ Acknowledgements
Appendix A Motivating Experiments
Additional details
For Figure˜2a, we use the same demonstrations used in our main experiments (see Appendix˜B for more details). For SAC and TD3 experiments with action sequences in Figure˜2b, we implemented our code based on the official HumanoidBench repository. We use the hyperparameters in the repository for training SAC agents. For TD3, we use the standard deviation of 0.2 for exploration. We report the average target Q-values recorded throughout experiments. For Figure˜2c, we train SAC and CQN agents with 6 original actions and 294 no-op actions with [-1, 1] action bounds and use an environment wrapper that slices out no-op actions.
Experiment with 2D Point-mass environment
To further motivate the use of action sequence for value learning, we train DQN agents (mnih2015human) on 2D Point-mass environment with discrete action spaces. In particular, we train a Raw agent that trains with the raw discrete action space consisting of single-step accelerations parameterized by 8 discrete headings (cardinal and discrete directions) and 1 magnitude level, resulting in 8 total actions. We compare this against a Sequence agent that trains with the discrete action space that consists of smooth 5-step acceleration sequences parameterized by cubic Bezier curves instead of single-step accelerations. Both agents operate on a 2D double-integrator environment where the goal is to reach a target position. In Figure˜9, as expected, we find that training DQN agent with pre-defined action sequences lead to faster convergence. Our main experimental results in Section˜4 further show that, CQN-AS can achieve similar benefit of using action sequences without pre-defined set of action sequences on various challenging continuous control benchmarks.
(a) 2D Point-mass task
(b) Experimental Results
Figure 9: 2D Point-mass experiments (a) We consider a simple toy environment where the goal is to reach the region around a randomly spawned goal point. (b) We show that, as expected, training with pre-defined action sequences lead to faster convergence. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 10 runs.